OpenStudy (nali):

plz help.. when 15 is divided by the positive integers k, the reminder is 3. for how many different values of k is this true? 1 2 3 4 5

OpenStudy (nali):

i am studying for the sat and i need to know the simpliest and fastest way to do this type of problem

OpenStudy (vijay):

2

OpenStudy (nali):

no it is 3 in my book

OpenStudy (vijay):

yes right 3

OpenStudy (vijay):

4 , 6 , 12

OpenStudy (nali):

can u plz explain in a more clear way?

OpenStudy (vijay):

if u divide 15 by 4 , u will get answer 3 & reminder 3.

jimthompson5910 (jim_thompson5910):

15 = nk + r 15 = nk + 3 15-3 = nk 12 = nk nk = 12 So any number that is a factor of 12 will work. However, 2 does NOT work since 15/2 = 7 remainder 1.

OpenStudy (vijay):

if u divide 15 by 6 answer 2 & reminder 3

OpenStudy (nali):

@jim_thompson5910 what is n ?

jimthompson5910 (jim_thompson5910):

n is any integer so if n = 2 and k = 6 for instance, then nk = 12 2*6 = 12 12 = 12

jimthompson5910 (jim_thompson5910):

To satisfy that equation, n has to be less than 12

OpenStudy (nali):

i dont get how u formed ur equation 15 = nk + r

jimthompson5910 (jim_thompson5910):

When you divide 15 by k, you get 15/k This will give you a quotient n and some remainder r So 15/k = n+r/k which can be rearranged to 15 = nk + r

OpenStudy (nali):

ok @jim_thompson5910 now i get it but how do u find the answer from this equation ?

jimthompson5910 (jim_thompson5910):

You're simply looking for numbers that are factors of 12 So just count the factors of 12. Exclude 1, 2, and 3 because they won't work here.

OpenStudy (nali):

ok so how did u know that u have to exclude those numbers

OpenStudy (nali):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

because 15/1 = 15 remainder 0 15/2 = 7 remainder 2 15/3 = 5 remainder 0

jimthompson5910 (jim_thompson5910):

but we want a remainder of 3

OpenStudy (nali):

OH ok thanku so much

jimthompson5910 (jim_thompson5910):

oops, made a typo 15/2 = 7 remainder 1

jimthompson5910 (jim_thompson5910):

but it still doesn't work

OpenStudy (nali):

ok